Find the flux of Earth's magnetic field of magnitu
Answer #1

here,

magnatic field, B = 5*10^-5 T

area, A = 30 cm^2 = 0.0030 m^2

From gauss law ::
rate of change of flux, Fi = B * A * CosA

Where, A = Angle b/w magnatic field and surface area

Part A:
Fi = B*A*Cos90
Fi = 0

Part b:
Fi = B*A*Cos40
Fi = 5*10^-5 * 0.0030 * Cos40
Fi = 1.149 * 10^-7 T.m^2

Part C:
Fi = B * A * Cos0
Fi = 5*10^-5 * 0.0030 * Cos0
Fi = 1.5 * 10^- 7 T.m^2

NOTE : flux will be max when A = 0 degrees and min when A = 90 degree